How many ways can a committee of 2 boys and 2 girls and an alternate of either gender be chosen from 5 boys and 6 girls?
Answer by Edwin McCravy(20054)

We can choose the two boys any of 5C2 or 10 ways.
For each of the 10 ways to choose the boys, we can
choose the two girls any of 6C2 or 15 ways.  

That's 10·15 or 150 ways to choose the two boys and 
the two girls.  

For each of those 150 ways which we can choose the two 
boys and the two girls, we can choose the alternate as 
any of the 7 remaining people.  That's 150·7 or 1050 
ways.

Answer: 5C2·6C2·7 = 10·15·7 = 1050 ways.

Edwin